# Switching the Loss Reduces the Cost in Offline Reinforcement Learning

### v2

@@ A.2 Concentration for the log-loss estimator
Fix a set $\mathcal{X}$, which, for the sake of avoiding measurability issues is assumed to be finite. Let $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$ independent, identically distributed random variables taking values in $\mathcal{X}\times[0,1]$. Let $f^{\star}$ be the regression function underlying $\nu$: $f^{\star}(x)=\mathbb{E}\,\left[Y_{1}\,|\,X_{1}=x\right]$. Let $\mathcal{F}\subset\{f\,:\,f:\mathcal{X}\to[0,1]\}$ be a finite set of $[0,1]$-valued functions with domain $\mathcal{X}$. Recall the log-loss estimator:
@@
$\ell_{\log}(y;\,y^{\prime}):=y^{\prime}\log\frac{1}{y}+(1-y^{\prime})\log\frac{1}{1-y}\,,$
where we define $00\log\infty=\lim_{x\to 0}x\log 1/x=$ 0. Foster and Krishnamurthy [2021] show the following concentration result for $\hat{f}_{\log}$, which we will need:
@@
Suppose $f^{\star}\in\mathcal{F}$. Let $D_{n}=((X_{1},Y_{1}),\dots,(X_{n},Y_{n}))$. Then, for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
@@
where $\nu$ denotes the common distribution over $X_{1},\dots,X_{n}$.
@@
The result follows from the last equation on page 24 of the arXiv version of the paper by Foster and Krishnamurthy [2021] with $A=1$. ∎
@@
In this section we give the main steps of the proof of Theorem 4.7. For the benefit of the reader, we first reproduce the text of the theorem:
@@
Given a dataset $D=\{(S_{i},A_{i},C_{i},S_{i}^{\prime})\}_{i=1}^{n}$ with $n\in\mathbb{N}$ and a finite function class $\mathcal{F}\subseteq(\mathcal{S}\times\mathcal{A}\rightarrow[0,1])$ that satisfy Assumptions 4.2, 4.3, 4.5 and 4.6, $B=1$, it holds with probability $1-\delta$, the output policy of FQI-log after $k$ iterations, $\pi_{k}=\pi_{f_{k}}$, satisfies
$\bar{v}^{\pi_{k}}-\bar{v}^{\star}\leq\tilde{C}\left(\frac{1}{(1-\gamma)^{2}}\sqrt{\frac{\bar{v}^{\star}C\log\left(\lvert\mathcal{F}\rvert^{2}/\delta\right)}{n}}+\frac{C\log\left(\lvert\mathcal{F}\rvert^{2}/\delta\right)}{(1-\gamma)^{4}n}+\frac{\gamma^{\frac{k}{2}}}{1-\gamma}+\frac{\gamma^{k}}{(1-\gamma)^{2}}\right)\,.$ (9)
where $\tilde{C}>0$ is an absolute constant.
@@
The first proposition shows that the error of a policy that is greedy with respect to an action-value function $f:\mathcal{S}\times\mathcal{A}\to[0,\infty)$ can be bounded by the triangular discrimination distance between the action-value function and $q^{*}$, the optimal action-value function in our MDP. To state this proposition, for $f$ as above we define $\Delta_{f}:\mathcal{S}\times\mathcal{A}\to[0,\infty)$, as the pointwise triangular deviation of $f$ from $q^{*}$:
@@
To state the proposition, recall that for a distribution $\eta$ over the states and for a stationary policy $\pi$, we let $\eta\times\pi$ denote the joint probability distribution over the state-action pairs resulting from first sampling a state $S$ from $\eta$, followed by choosing an action $A$ at random from $\pi(S)$. size=,color=red!20!white,]Cs: review notation/terminology for policies. also, this is just applying a Markov kernel.. Same as $\nu P$. So why the $\times$? With this, the first proposition is as follows:
@@
When the above proposition is applied to $f=f_{k}$, the action-value function obtained in the $k$th iteration of our algorithm, we see that it remains to bound $D_{f_{k}}$. The bound will be based on the second proposition:
@@
For any admissible distribution $\nu$ over $\mathcal{S}\times\mathcal{A}$ that may also depend on the data $D_{n}$,size=,color=red!20!white,]Cs: I commented out the assumptions.. Usually one says before the lemmas that the assumptions hold.. for any $10<\delta<$ 1, $k\geq 1$, with probability $1-\delta$,
@@
The proof of this proposition uses (i) showing that $\mathcal{T}$ enjoys some contraction properties with respect to appropriately chosen Hellinger distances; (ii) using these contraction properties to show that the Hellinger distance between $f_{k}$ and $q^{*}$ is controlled by the Hellinger distances between $f_{k}$ and $\mathcal{T}f_{k}$, and then using the results of the previous section to show that these are controlled by the algorithm.
@@
also holds, on the same event when Eq. 13 holds. Plugging these bounds into Eq. 11, we get that with probability at least $1-\delta$,
@@
The analysis in this section is inspired by the proof of Lemma 1 of Foster and Krishnamurthy [2021]. We deviate from their analysis to avoid introducing an extra $|\mathcal{A}|$ factor in the bounds.
@@
Lemma B.4 (Performance Difference Lemma of Kakade and Langford).
@@
See Lemma 6.1 by Kakade and Langford [2002]. ∎
The next lemma upper bounds the one-norm distance between a nonnegative-valued function $f:\mathcal{S}\times\mathcal{A}\to[0,\infty)$ and $q^{*}$ in terms of appropriate norms of $\Delta_{f}$ and $\xi_{f}$.
@@
Combining this with Eq. 20 gives the desired inequality.
@@
B.2 Bounding the triangular deviation between $f_{k}$ and $q^{*}$: Proof of Proposition B.3
As explained earlier, the analysis in this section uses contraction arguments that have a long history in the analysis of dynamic programming algorithms in the context of MDPs. The novelty is that we need to bound the triangular deviation to $q^{*}$. As this has been shown to be upper bounded by the Hellinger distance (Corollary A.2), we switch to Hellinger distances and establishes contraction properties of $\mathcal{T}$ with respect to such distances. This required new proofs. The change of measure arguments used in the “error propagation analysis” are standard.
@@
$\leq\sum_{s\in S}\eta(s)\left(\sqrt{f(s,\pi_{f,g}(s))}-\sqrt{g(s,\pi_{f,g}(s))}\right)^{2}$ (by Eq. 24)
@@
size=,color=red!20!white,]Cs: not anymore it seems! change back.. let’s discuss..? I wrote $\mathcal{T}f=c+\gamma Pf^{\wedge}$ on purpose.. With the above, you also need to say, operator application takes higher precedence than function evaluation. Which is of course how things must be, but still, need to mention these conventions. And from the above the vector/matrix notation is gone. But that view is useful below. Only change things you are absolutely sure about.. I don’t want to go over things again (sorry, just trying to save time). size=,color=green!20!white,]AA: No worries, i changed it back size=,color=red!20!white,]Cs: Note that I like to use $S$, $A$ for cardinalities; keeping $\mathcal{S}$ and $\mathcal{A}$ for the sets. This requires work.
@@
$\left\|\sqrt{f_{k}}-\sqrt{q^{*}}\right\|_{2,\nu}\leq\gamma^{\frac{k}{2}}+\frac{2\sqrt{C}}{1-\gamma}\max_{1\leq\tau\leq k}\left\|\sqrt{f_{\tau}}-\sqrt{\mathcal{T}f_{\tau-1}}\right\|_{2,\mu}\,.$
@@
Define $(\nu_{i})_{0\leq i\leq k}$ via $\nu_{k}=\nu$ and for $0\leq i<k$, let $\nu_{i}=(\nu_{i+1}P)\times\pi_{f_{i},q^{*}}$. Note that by assumption, $\nu_{k}$ is admissible. It then follows that $\nu_{i}$ for $0\leq i<k$ is also admissible. Indeed, if for some $0\leq i<k$, $\pi=(\pi_{0},\pi_{1},\dots)$ is the nonstationary policy that realizes $\nu_{i+1}$ in step $s\geq 0$, $\pi^{\prime}=(\pi_{0},\pi_{1},\dots,\pi_{s},\pi_{f_{i},q^{*}},\pi_{s+1},\dots)$ is a policy that realizes $\nu_{i}$ in step $s+1$.
@@
$\left\|\sqrt{f_{k}}-\sqrt{q^{*}}\right\|_{2,\nu}$ $\leq\sqrt{C}\,\sum_{t=1}^{k}\gamma^{\frac{{k-t}}{2}}\,\left\|\sqrt{f_{t}}-\sqrt{\mathcal{T}f_{t-1}}\right\|_{2,\mu}+\gamma^{\frac{k}{2}}\left\|\sqrt{f_{0}}-\sqrt{q^{\star}}\right\|_{2,\nu_{0}}$
$\leq\sqrt{C}\,\underbrace{\sum_{t=1}^{k}\gamma^{\frac{{k-t}}{2}}\,\left\|\sqrt{f_{t}}-\sqrt{\mathcal{T}f_{t-1}}\right\|_{2,\mu}}_{S:=}+\gamma^{\frac{k}{2}}\,,$
where the second inequality holds because by assumption, $f_{0}$ takes values in $[0,1]$ and so does $q^{*}$. Hence, $(\sqrt{f_{0}}-\sqrt{q^{\star}})^{2}\leq 1$ and thus $\left\|\sqrt{f_{0}}-\sqrt{q^{\star}}\right\|_{2,\nu_{0}}\leq 1$.
@@
$S$ $\leq\max_{\tau\in[1,\dots,k]}\left\|\sqrt{f_{\tau}}-\sqrt{\mathcal{T}f_{\tau-1}}\right\|_{2,\mu}\sum_{t=1}^{k}\gamma^{\frac{k-t}{2}}$
@@
For $g:\mathcal{S}\times\mathcal{A}\to[0,1]$, let $\hat{f}_{g}$ be the function learned by regressing on $g$ via log-loss, i.e., size=,color=red!20!white,]Cs: perhaps better to introduce $\hat{\mathcal{T}}$ so that $\hat{\mathcal{T}}g=\hat{f}_{g}$ (and not introduce $\hat{f}_{g}$.. “empirical operator”.. more expressive of what is going on..
@@
size=,color=red!20!white,]Cs: The last inequality when $\tau=1$ requires that $f_{0}\in\mathcal{F}$ and at some point we choose $f_{0}=0$. So either need a new assumption, or drop the choice $f_{0}=0$! Since this applies for any $1\leq\tau\leq k$, all that remains is to bound the right-hand side of the last display. We will use Theorem A.3 for this purpose. This result can be applied because, on the one hand, by Assumption 4.2, $\mathbb{E}\,[C_{i}+\gamma g^{\wedge}(S_{i}^{\prime})|S_{i},A_{i}]=\mathcal{T}g\,(S_{i},A_{i})$, and by Assumption 4.5, $\mathcal{T}g\in\mathcal{F}$ whenever $g\in\mathcal{F}$ and because, again, by Assumption 4.2, $(S_{i},A_{i},C_{i},S_{i+1}^{\prime})$ are independent, identically distributed random variables for $i=1,\dots,n$. size=,color=red!20!white,inline]Cs: A contentious issue is the assumption that $\mathcal{F}$ is finite and yet $\mathcal{T}\mathcal{F}\subset\mathcal{F}$. Will this ever be satisfied? Take the finite case for example. Already in this case, this would require that $\mathcal{T}^{i}0\in\mathcal{F}$ for $i=0,1,\dots$ assuming that $0\in\mathcal{F}$.. So this will need to be worked on. Thus, Theorem A.3 together with a union bound and recalling that the distribution of $(S_{i},A_{i})$ is $\mu$ gives that for any $10<\delta<$ 1,

### v5

@@ A.2 Concentration for the log-loss estimator
Fix a set $\mathcal{X}$, which, for the sake of avoiding measurability issues, is assumed to be finite. Let $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$ be independent, identically distributed random elements taking values in $\mathcal{X}\times[0,1]$. Let $f^{\star}$ be the regression function underlying $\nu$: $f^{\star}(x)=\mathbb{E}\,\left[Y_{1}\,|\,X_{1}=x\right]$. size=,color=orange!20!white,]SR: should $\nu$ be defined before this, or is this meant to serve as definition? Let $\mathcal{F}\subseteq[0,1]^{\mathcal{X}}$ be a finite set of $[0,1]$-valued functions with domain $\mathcal{X}$. Recall the log-loss estimator: size=,color=orange!20!white,]SR: first time this is defined
@@
$\ell_{\log}(y;\,y^{\prime})=y^{\prime}\log\frac{1}{y}+(1-y^{\prime})\log\frac{1}{1-y}\,,$
where we define $00\log\infty=\lim_{x\to 0}x\log 1/x=$ 0. Foster & Krishnamurthy (2021) show the following concentration result for $\hat{f}_{\log}$, which we will need:
@@
Suppose $f^{\star}\in\mathcal{F}$. Let $D_{n}=\{(X_{i},Y_{i})\}_{i=1}^{n}$. Then, for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
@@
where $\nu$ denotes the common distribution of $X_{1},\dots,X_{n}$. size=,color=orange!20!white,]SR: does this use of $\nu$ conflict with the one just above? I thought if $f^{\star}$ was the regression function of $\nu$ then it should be the dist of $(X_{i},Y_{i})$?
@@
The result follows from the last equation on page 24 of the arXiv version of the paper by Foster & Krishnamurthy (2021) with $A=1$. ∎
@@
In this section we give the main steps of the proof of Theorem 5.1. For the benefit of the reader, we first reproduce the text of the theorem.
@@
Given a dataset $D_{n}=\{(S_{i},A_{i},C_{i},S_{i}^{\prime})\}_{i=1}^{n}$ with $n\in\mathbb{N}$ and a finite function class $\mathcal{F}\subseteq[0,1]^{\mathcal{S}\times\mathcal{A}}$ that satisfy Assumptions 3.1, 3.2 and 3.4, it holds with probability $1-\delta$ that the output policy of FQI-log after $k$ iterations, $\pi_{k}=\pi_{f_{k}}$, satisfies
$\bar{v}^{\pi_{k}}-\bar{v}^{\star}\leq\tilde{C}\left(\frac{1}{(1-\gamma)^{2}}\sqrt{\frac{\bar{v}^{\star}C\log\left(\lvert\mathcal{F}\rvert^{2}/\delta\right)}{n}}+\frac{C\log\left(\lvert\mathcal{F}\rvert^{2}/\delta\right)}{(1-\gamma)^{4}n}+\frac{\gamma^{\frac{k}{2}}}{1-\gamma}+\frac{\gamma^{k}}{(1-\gamma)^{2}}\right)\,.$
where $\tilde{C}>0$ is an absolute constant. size=,color=orange!20!white,]SR: since we’re changing the statement to have the full equations, I think we should also add in names of assumptions (reasonable distribution, concentrability with coefficient $C$, and completeness)
@@
The first proposition shows that the error of a policy that is greedy with respect to an action-value function $f:\mathcal{S}\times\mathcal{A}\to[0,\infty)$ can be bounded by the triangular discrimination between the action-value function and $q^{\star}$, the optimal action-value function in our MDP. To state this proposition, for $f$ as above, we define $\Delta_{f}:\mathcal{S}\times\mathcal{A}\to[0,\infty)$, the pointwise triangular deviation of $f$ from $q^{\star}$: size=,color=orange!20!white,]SR: above it’s called discrimination but now it’s called deviation. Can we call it discrimination everywhere?
@@
To state the proposition, recall that for a distribution $\eta$ over the states and a stationary policy $\pi$, we let $\eta\times\pi$ denote the joint probability distribution over the state-action pairs resulting from first sampling a state $S\sim\eta$ and then an action $A\sim\pi(S)$. size=,color=red!20!white,]Cs: review notation/terminology for policies. also, this is just applying a Markov kernel.. Same as $\nu P$. So why the $\times$? size=,color=orange!20!white,]SR: we’ll have to introduce this above if we’re gonna say recall With this, the first proposition is as follows:
@@
When the above proposition is applied to $f=f_{k}$, the action-value function obtained in the $k^{\text{th}}$ iteration of our algorithm, we see that it remains to bound $D_{f_{k}}$. The bound will be based on the second proposition:
@@
For any admissible distribution $\nu$ over $\mathcal{S}\times\mathcal{A}$ that may also depend on the data $D_{n}$,size=,color=red!20!white,]Cs: I commented out the assumptions.. Usually one says before the lemmas that the assumptions hold.. for any $\delta\in(0,1)$, $k\geq 1$, with probability $1-\delta$,
@@
The proof of this proposition uses (i) showing that $\mathcal{T}$ enjoys some contraction properties with respect to appropriately chosen Hellinger distances; (ii) using these contraction properties to show that the Hellinger distance between $f_{k}$ and $q^{\star}$ is controlled by the Hellinger distances between $f_{k}$ and $\mathcal{T}f_{k}$, and then using the results of the previous section to show that these are controlled by the algorithm.
@@
also holds, on the same event when Equation 12 holds. Plugging these bounds into Equation 10, we get that with probability at least $1-\delta$,
@@
The analysis in this section is inspired by the proof of Lemma 1 of Foster & Krishnamurthy (2021). We deviate from their analysis to avoid introducing an extra $|\mathcal{A}|$ factor in the bounds.
@@
Lemma B.4 (Performance Difference Lemma of Kakade & Langford).
@@
See Lemma 6.1 by Kakade & Langford (2002). ∎
The next lemma upper bounds the one-norm distance between a nonnegative-valued function $f:\mathcal{S}\times\mathcal{A}\to[0,\infty)$ and $q^{\star}$ in terms of appropriate norms of $\Delta_{f}$ and $\xi_{f}$.
@@
Combining this with Equation 19 gives the desired inequality.
@@
B.2 Bounding the triangular deviation between $f_{k}$ and $q^{\star}$: Proof of Proposition B.3
As explained earlier, the analysis in this section uses contraction arguments that have a long history in the analysis of dynamic programming algorithms in the context of MDPs. The novelty is that we need to bound the triangular deviation to $q^{\star}$. As this has been shown to be upper bounded by the Hellinger distance (Corollary A.2), we switch to Hellinger distances and establishes contraction properties of $\mathcal{T}$ with respect to such distances. This required new proofs. The change of measure arguments used in the “error propagation analysis” are standard.
@@
$\leq\sum_{s\in S}\eta(s)\left(\sqrt{f(s,\pi_{f,g}(s))}-\sqrt{g(s,\pi_{f,g}(s))}\right)^{2}$ (by Equation 23)
@@
size=,color=red!20!white,]Cs: not anymore it seems! change back.. let’s discuss..? I wrote $\mathcal{T}f=c+\gamma Pf^{\wedge}$ on purpose.. With the above, you also need to say, operator application takes higher precedence than function evaluation. Which is of course how things must be, but still, need to mention these conventions. And from the above the vector/matrix notation is gone. But that view is useful below. Only change things you are absolutely sure about.. I don’t want to go over things again (sorry, just trying to save time).
@@
$\left\|\sqrt{f_{k}}-\sqrt{q^{\star}}\right\|_{2,\nu}\leq\gamma^{\frac{k}{2}}+\frac{2\sqrt{C}}{1-\gamma}\max_{1\leq\tau\leq k}\left\|\sqrt{f_{\tau}}-\sqrt{\mathcal{T}f_{\tau-1}}\right\|_{2,\mu}\,.$
@@
Define $(\nu_{i})_{0\leq i\leq k}$ via $\nu_{k}=\nu$ and for $0\leq i<k$, let $\nu_{i}=(\nu_{i+1}P)\times\pi_{f_{i},q^{\star}}$. Note that by assumption, $\nu_{k}$ is admissible. It then follows that $\nu_{i}$ for $0\leq i<k$ is also admissible. Indeed, if for some $0\leq i<k$, $\pi=(\pi_{0},\pi_{1},\dots)$ is the nonstationary policy that realizes $\nu_{i+1}$ in step $s\geq 0$, $\pi^{\prime}=(\pi_{0},\pi_{1},\dots,\pi_{s},\pi_{f_{i},q^{\star}},\pi_{s+1},\dots)$ is a policy that realizes $\nu_{i}$ in step $s+1$.
@@
$\left\|\sqrt{f_{k}}-\sqrt{q^{\star}}\right\|_{2,\nu}$ $\leq\sqrt{C}\,\sum_{j=1}^{k}\gamma^{\frac{{k-j}}{2}}\,\left\|\sqrt{f_{j}}-\sqrt{\mathcal{T}f_{j-1}}\right\|_{2,\mu}+\gamma^{\frac{k}{2}}\left\|\sqrt{f_{0}}-\sqrt{q^{\star}}\right\|_{2,\nu_{0}}$
$\leq\sqrt{C}\,\underbrace{\sum_{j=1}^{k}\gamma^{\frac{{k-j}}{2}}\,\left\|\sqrt{f_{j}}-\sqrt{\mathcal{T}f_{j-1}}\right\|_{2,\mu}}_{S:=}+\gamma^{\frac{k}{2}}\,,$
where the second inequality holds because by assumption, $f_{0}$ takes values in $[0,1]$ and so does $q^{\star}$. Hence, $(\sqrt{f_{0}}-\sqrt{q^{\star}})^{2}\leq 1$ and thus $\left\|\sqrt{f_{0}}-\sqrt{q^{\star}}\right\|_{2,\nu_{0}}\leq 1$.
@@
$S$ $\leq\max_{\tau\in[1,\dots,k]}\left\|\sqrt{f_{\tau}}-\sqrt{\mathcal{T}f_{\tau-1}}\right\|_{2,\mu}\sum_{j=1}^{k}\gamma^{\frac{k-j}{2}}$
@@
For $g:\mathcal{S}\times\mathcal{A}\to[0,1]$, let $\hat{f}_{g}$ be the function learned by regressing on $g$ via log-loss, i.e.,
@@
Since this applies for any $1\leq\tau\leq k$, all that remains is to bound the right-hand side of the last display. We will use Theorem A.3 for this purpose. This result can be applied because, on the one hand, by Assumption 3.1, $\mathbb{E}\,[C_{i}+\gamma g^{\wedge}(S_{i}^{\prime})|S_{i},A_{i}]=\mathcal{T}g\,(S_{i},A_{i})$, and by Assumption 3.4, $\mathcal{T}g\in\mathcal{F}$ whenever $g\in\mathcal{F}$ and because, again, by Assumption 3.1, $(S_{i},A_{i},C_{i},S_{i+1}^{\prime})$ are independent, identically distributed random variables for $i=1,\dots,n$. size=,color=red!20!white,inline]Cs: A contentious issue is the assumption that $\mathcal{F}$ is finite and yet $\mathcal{T}\mathcal{F}\subset\mathcal{F}$. Will this ever be satisfied? Take the finite case for example. Already in this case, this would require that $\mathcal{T}^{i}0\in\mathcal{F}$ for $i=0,1,\dots$ assuming that $0\in\mathcal{F}$.. So this will need to be worked on. Thus, Theorem A.3 together with a union bound and recalling that the distribution of $(S_{i},A_{i})$ is $\mu$ gives that for any $10<\delta<$ 1,